A new computer can complete a complex biochemistry calculation in 7 hours less than an older-model computer. Working together, the new computer and the old computer can complete the calculation in 17 hours. How long would it take the new computer to complete the calculation by itself? Round to the nearest tenth of an hour.

Answer by jorel1380(3719)

Let n and n-7 be the old and new computers, respectively. Then:
1/n + 1/n-7=1/17
17(n-7)+17n=n^2-7n
n^2-41n+119=0
n=37.85655485770978
n-7=30.85655495770978 hours neede by the new computer, alone.🙂🙂🙂🙂